Marshall has received an inheritance and wants to invest a sum of money today that will yield $4,800 at the end of each of the next 10 years. Assuming he can earn an interest rate of 5% compounded annually, how much of his inheritance must he invest today? (PV of $1, FV of $1, PVA of $1, and FVA of $1) (Use appropriate factor(s) from the tables provided.)
a. 96,000.00
b. 37,064.16
c. 48,000.00
d. 45,600.00
e. 43,320.00

Portia Grant is an employee who is paid monthly. For the month of January of the current year, she earned a total of 8,738. The FICA tax for social security is 6.2% of the first $118,500 of employee earnings each calendar year and the FICA tax rate for Medicare is 1.45% of all earnings. The FUTA tax rate of 0.6% and the SUTA tax rate of 5.4% are applied to the first $7,000 of an employee's pay. The amount of federal income tax withheld from her earnings was $1,449.87. Her net pay for the month is: (Round your intermediate calculations to two decimal places.)
a. 6,566.00
b. 7,161.43
c. 5,116.13
d. 6,185.67
e. 6,619.67

On January 1, Parson Freight Company issues 7.5%, 10-year bonds with a par value of $2,600,000. The bonds pay interest semiannually. The market rate of interest is 8.5% and the bond selling price was $2,423,327. The bond issuance should be recorded as:
a. Debit Cash $2,600,000; credit Bonds Payable $2,600,000.
b. Debit Cash $2,423,327; credit Bonds Payable $2,423,327.
c. Debit Cash $2,423,327; debit Interest Expense $176,673; credit Bonds Payable $2,600,000.
d. Debit Cash $2,423,327; debit Discount on Bonds Payable $176,673; credit Bonds Payable $2,600,000.
e. Debit Cash $2,600,000; credit Bonds Payable $2,423,327; credit Discount on Bonds Payable $176,673.

Which column (i) and row (n) would you use from a present value or future value table for 10% interest compounded quarterly for 6 years?
a. (i) = 2.50%, (n) = 24
b. (i) = 2.50%, (n) = 10
c. (i) = 5%, (n) = 12
d. (i) = 10%, (n) = 6
e. (i) = 5%, (n) = 24
Answer #1
1) Marshall:
The amount to be deposited today is the PV of the annuity of $4800
= 4800*PVIFA(5,10) = 4800*7.7217 = 37064.16
Answer: Option [b] $37,064.16
2) Portia Grant:
Gross earnings 8738.00
Deductions:
FICA tax at 6.2% on 8738 541.76
Medicare at 1.45% on 8738 126.70
Income tax withheld 1449.87 2118.33
Net pay for the month 6619.67
Answer: Option [e] $6,619.67
FUTA and SUTA are payable by the employer.
3) Parson Freight company:
Answer: Option [d] Debit Cash $2,423,327; debit Discount on Bonds Payable $176,673; credit Bonds Payable $2,600,000.
4) Column to be used for PF or FV: a. (i) = 2.50%, (n) = 24
i = 10/4 = 2.5%
n = 6*4 =24